Find the value of b for which the equation
( ye^(2xy)+x)dx+bxe^(2xy)dy= 0
is exact, and hence solve it for that value of b
1
Expert's answer
2020-03-27T14:21:22-0400

The given equation is "(ye^{2xy}+x)dx+bxe^{2xy}dy=0" .

We know that a differential equation "Mdx+Ndy=0" is exact if


"\frac{\partial M}{\partial y}=\frac{\partial N}{\partial x}"

Therefore, according to question,

"M=(ye^{2xy}+x) \ and \ N=bxe^{2xy}"

Now,

"\frac{\partial (ye^{2xy}+x)}{\partial y}=e^{2xy}+2xye^{2xy}=(1+2xy)e^{2xy}.......(1)"

and

"\frac{\partial (bxe^{2xy})}{\partial x}=be^{2xy}+2xybe^{2xy}=(1+2xy)be^{2xy}........(2)"

from equation (1) and (2) ,we have the given differential equation is exact if "b=1" .

Now , putting b=1 in the given differential equation we get

"(ye^{2xy}+x)dx+xe^{2xy}dy=0"

"\implies \ ye^{2xy}dx+xdx+xe^{2xy}dy=0"

"\implies \ e^{2xy}(ydx+xdy)+xdx=0"

"\implies \ e^{2xy}d(xy)+xdx=0"

Integrating, we get

"\frac{1}{2}e^{2xy}+\frac{x^2}{2}=c"

where "c" is an integration constant.